Given the half-cells:

Ni/Ni+2 and Fe/Fe+2

Chose the anode and cathode. Calculate and ΔG° for this reaction. Calculate what the new cell potential (E) would be if the concentrations were [Ni+2] = 0.40 M and [Fe+2] = 0.30 M. Calculate how long you would have to run the battery at 6 amps to make the E drop by 0.03 volts, and calculate how many grams of anode you lose by that time.

Using the equations:

ΔG° = -nFE°

ln K = (nE°) / 0.0257

E = E° - (0.0257 / n) ln Q

Charge (coulombs) = Amps x time (seconds)
Answer #1

The half cell reduction potentials are :

Ni/Ni+2 = -0.257 Volts

Fe/Fe+2 = -0.447 Volts

Answer 1:

According to the above data: Anode is Fe/Fe2+ half cell and Cathode is Ni/Ni2+ half cell. Oxidation takes place on anode and reduction takes place in cathode.

reaction is represented as=

Fe + Ni2+ ---------------> Fe2+ + Ni

Answer 2:

Now, Eocell = Ecathode - Eanode

E0cell= (-0.257)-(-0.447) = 0.19 Volts

Answer 3:

ΔG°reaction = -nFE0cell ,

where n (number of electron change in reaction) = 2 mol e-, F (Faraday's constant)= 96485 coulomb/mol e-

ΔG°reaction = -2*96485*(0.19) = 36664.3 coulomb.volt = 36664.3 Joule = 36.664 kJ

Answer 4:

Ecell = E0cell - (0.0591/n)*log(Q)

where Q= [Fe2+]/[Ni2+]

We have to find the E value when the concentrations are ([Ni2+]=0.40M and [Fe2+]=0.30M

Ecell = 0.19 - (0.0591/2)*log(0.40/0.30) = 0.19 - (0.02955)*.124938 = 0.18630 Volt

Ecell = 0.18630 Volt